MOM Report Victoria Blackmore VC-141 16/06/11 1/7 Hall Update        Substation upgrade completed.  Power to equipment back on. Water restored to hall  Turned on on 07/06/11.  Conventional.

Report
MOM Report
Victoria Blackmore
VC-141
16/06/11
1/7
Hall Update







Substation upgrade completed.
 Power to equipment back on.
Water restored to hall
 Turned on on 07/06/11.
 Conventional magnets tested – all OK bar D1 maybe (see later).
Decay Solenoid cooled down.
EMR and TOF arrived on Monday.
 Lots of busy people in the hall installing detectors!!
Control room side sign-in sheet now next to hall entrance.
 Plus handy hat storage!
Weekly work list posted on MLCR door (Andy/Vishal).
ISIS now in shutdown period (until June 26th).
2/7
Systems Test


Systems test carried out on 13/06/11
 Question: What is working after the long shutdown and power
upgrade?
 Answer: Ask again later…
Long list of problems.. Some are fixed/being fixed now.
 Tweaks needed to the alarm handler (DS flow rates).
 Couldn’t turn on CAEN 127 HV remotely.



Had to do this in person, would be much better if it could be done
from the MLCR.
DS power supply blew a capacitor, a bit of smoke and blackened
fingers followed during investigation – fixed!
Run walkthrough checklist specifies the wrong keys for getting
the beamstop keys

Not major.. Just confusing.
3/7
Systems Test

Problems cont…
 PPS:








LH2 mezzanine door latch upside-down.
RF cage (west) door misaligned – fixed!
Grey door next to RF cage needs a padlock (11kV inside!) – fixed.
Main door bob requires excessive force to reset – in progress.
Inner main (personnel) door can be rattled, dropping hall search
condition.
Key to inner main door must be aligned correctly, or else you can
lock yourself inside the hall during a controlled access!
Need a permanent intercom by the personnel door.
Good news: Controlled access procedure worked well!
4/7
Systems Test

Problems cont…
 Target:




Oscilloscope wouldn’t turn on (lit black screen only). Swapped for
another oscilloscope, which did work.
Couldn’t get machine trigger, so no dipping since…
DAQ trigger kept tripping its power supply (4A crate on a 2A
fuse!).
Conventional magnets:





All OK except (maybe) D1
When ramping to high current (e.g. ~300A for reference run settings)
temperature gradient across magnet is much higher than others.
Usually expect 1-2C, D1 was around 15-20C.
Lowering the current returned this to ‘normal’.
Vishal has ordered some temperature sensors that may help
diagnose the problem…
5/7
Near Miss!


Wearing hard hats and safety boots in the
MICE hall is vital.
A fluorescent tube fell from the hall ceiling
last night onto the DSA area.



No-one hurt, no damage done.
Lots of shattered glass.
Lucky!
6/7
Summary

Busy time in the hall right now



Fixing ‘bugs’
Installing EMR, TOF1, CKOVa
Need to finalise run shifts and gather shifters!
7/7

similar documents